### v1

# GaussianOcc: Fully Self-supervised and Efficient 3D Occupancy Estimation with Gaussian Splatting

Abstract
We introduce GaussianOcc, a systematic method that investigates the two usages of Gaussian splatting for fully self-supervised and efficient 3D occupancy estimation in surround views. First, traditional methods for self-supervised 3D occupancy estimation still require ground truth 6D poses from sensors during training. To address this limitation, we propose Gaussian Splatting for Projection (GSP) module to provide accurate scale information for fully self-supervised training from adjacent view projection. Additionally, existing methods rely on volume rendering for final 3D voxel representation learning using 2D signals (depth maps, semantic maps), which is both time-consuming and less effective. We propose Gaussian Splatting from Voxel space (GSV) to leverage the fast rendering properties of Gaussian splatting. As a result, the proposed GaussianOcc method enables fully self-supervised (no ground truth pose) 3D occupancy estimation in competitive performance with low computational cost (2.7 times faster in training and 5 times faster in rendering). The relevant code will be available in https://github.com/GANWANSHUI/GaussianOcc.git.
Figure 1: Problem setting of GaussianOcc. Given the surround image sequence and its corresponding 2D semantic annotation, GaussianOcc is able to perform 3D occupancy estimation without the need for the ground truth occupancy label and ground truth 6D pose for training.
1 Introduction
In recent years, both vision-centric and grid-centric perception approaches have gained significant attention in the autonomous driving industry and academia [35, 39]. Among these, surround view 3D occupancy estimation [32, 45, 51, 48] has emerged as a core perception task and a promising alternative to bird’s-eye view (BEV) methods [9, 29, 30]. To facilitate 3D occupancy estimation, several benchmarks have been developed for supervised training [45, 43, 41, 42], though these require substantial effort in 3D annotation. To reduce the burden of annotation, self-supervised [12, 54, 24, 5, 18] and weakly-supervised [37] learning approaches based on volume rendering have been proposed [36, 46, 11]. Volume rendering allows 3D representation learning using 2D supervision signals, such as 2D semantic maps and depth maps, thereby eliminating the need for extensive 3D annotation during training.
The existing methods [54, 24] achieve self-supervised learning through volume rendering, where the 2D semantic map supervision is derived from open-vocabulary semantic segmentation [55], and the depth map supervision is obtained from self-supervised depth estimation [13]. However, these approaches face two significant limitations. First, volume rendering is performed at real-world scale, which requires the availability of ground truth 6D pose to calculate the multi-view photometric loss across sequential images. Second, the inefficiency in volume rendering poses a challenge, the same as in novel view synthesis tasks [27, 1, 19], due to the dense sampling operation required. These limitations impede the development of a more general and efficient paradigm for self-supervised 3D occupancy estimation.
To address the aforementioned limitations, we explore a fully self-supervised and efficient approach to 3D occupancy estimation using Gaussian splatting [27, 1]. Specifically, we introduce the use of Gaussian splatting to perform cross-view splatting, where the rendered image constructs a cross-view loss that provides scale information during joint training with the 6D pose network. This eliminates the need for a ground truth 6D pose during training. To improve rendering efficiency, we move away from the dense sampling required in traditional volume rendering. Instead, we propose performing Gaussian splatting directly from the 3D voxel space. In this approach, each vertex in the voxel grid is treated as a 3D Gaussian, and we directly optimize the attributes of these Gaussians—such as semantic features and opacity—within the voxel space. Through this novel approach, our proposed method, GaussianOcc, makes a progress toward fully self-supervised and efficient 3D occupancy estimation, as outlined in Figure 1.
The main contributions of this work are summarized as follows.
  • We introduce the first fully self-supervised method for efficient surrounding-view 3D occupancy estimation, featuring the exploration of the Gaussian splatting.
  • We propose Gaussian splatting for cross-view projection, which could provide the scale information to get rid of the need of a ground truth 6D pose during the training.
  • We propose Gaussian splatting from voxel space achieving competitive performance with 2.7 times faster training and 5 times faster rendering compared with the volume rendering.
2 Related work
2.1 Surround view depth estimation
The surround view setting offers an ego-centric 360-degree perception solution [35, 39, 3, 14]. [17] introduces the surround view benchmark in the supervised setting, which learns the depth scale directly from the ground truth depth map. FSM [16] is a pioneering work in scale-aware surround view depth estimation relying on stereo constraint [53]. However, subsequent studies [44, 28] have found that reproducing the performance of [16] is challenging. Surrounddepth [44] improves the scale supervision signal using sparse point clouds from Structure-from-Motion (SFM). Building on the spatial and temporal constraints in FSM, [28] introduces a volume feature fusion module to enhance performance. For better performance, [38] proposes the temporal offline refinement strategy based on the multiple cameras and monocular depth refinement. These works [16, 44, 28, 38] all use traditional projection and index interpolation for cross-view synthesis, computing the loss between the synthesized view and target images. In contrast, our approach employs Gaussian splatting with a dedicated design, achieving better performance compared to the solely spatial constraint.
2.2 Surround view 3D occupancy estimation
Surround view 3D occupancy estimation has gained significant attention in recent years, with several benchmarks based on the Nuscene dataset [45, 41, 43]. In addition to the advanced architectures being proposed [32, 56, 51, 31], another research trend involves utilizing volume rendering for 3D supervision with 2D supervision [12, 37, 54, 24]. SimpleOcc [12] pioneered the use of volume rendering for 3D occupancy estimation, exploring both supervised and self-supervised learning. RenderOcc [37] extends semantic information for rendering. OccNeRF [54] and SelfOcc [24] share a similar approach by using 2D open-vocabulary semantic models to generate semantic maps for supervision. However, since volume rendering processes are conducted at real-world scale, these self-supervised methods [12, 54, 24] require ground truth 6D poses from sensors to provide the real-world scale for training. In contrast, we are exploring a solution that utilizes the overlap region in adjacent cameras to learn the real-world scale, eliminating the need for ground truth 6D poses.
2.3 3D Gaussian splatting
3D Gaussian splatting has become a popular method for modeling 3D and 4D scenes using well-posed images [27, 26, 8, 47], which has the property of fast rendering compared to the volume rendering in neural radiance field [36, 19, 11]. In the context of driving scenes, a line of research has focused on scene-specific reconstruction [22, 58, 49]. Our work, however, investigates the function of Gaussian splatting in a generalized setting. In this setting, existing methods generally construct 3D Gaussians from the unprojection of learned 2D Gaussian attributes [7, 57, 33, 40]. We also employ this unprojection approach, but uniquely, our approach constructs cross-view information from adjacent views to learn scale information through Gaussian splatting projection.
Additionally, two recent works, GaussianFormer [25] and GaussianBeV [6], are closely related to ours in their focus on 3D occupancy estimation and BEV prediction. However, our exploration diverges by focusing on two new properties that Gaussian splatting can contribute to occupancy estimation: scale-aware training and faster rendering.
3 Method
Figure 2: The overview of the proposed GaussianOcc. With a sequence of surround images, we employ a U-Net architecture to predict Gaussian attributes in the 2D image grid space for cross-view Gaussian splatting. This approach provides scale information in the joint training with the 6D pose network (Stage 1). For 3D occupancy estimation, we lift the 2D features to a 3D voxel space and propose voxel grid Gaussian splatting for fast rendering (Stage 2). We omit the line to from the 6D pose net to the loss in stage 1 for clarity.
3.1 Preliminaries
In this paper, we investigate the use of 3D Gaussian splatting [27] for fully self-supervised 3D occupancy estimation. 3D Gaussian Splatting (3D-GS) is a technique for modeling static 3D scenes using point primitives. Each point is represented as a scaled Gaussian distribution characterized by a 3D covariance matrix $\Sigma$ and a mean $\mu$. The Gaussian distribution for a point $\mathcal{X}$ is given by:
$G(\mathcal{X})=e^{-\frac{1}{2}(\mathcal{X}-\mu)^{T}\Sigma^{-1}(\mathcal{X}-\mu)}$ (1)
To enable efficient optimization via gradient descent, the covariance matrix $\Sigma$ is decomposed into a scaling matrix $\mathbf{S}$ and a rotation matrix $\mathbf{R}$ as follows:
$\Sigma=\mathbf{RSS}^{T}\mathbf{R}^{T}$ (2)
The projection of Gaussians from 3D space onto a 2D image plane involves a view transformation $\mathbf{W}$ and the Jacobian of the affine approximation of the projective transformation $\mathbf{J}$. The 2D covariance matrix $\Sigma^{\prime}$ is computed as:
$\Sigma^{\prime}=\mathbf{JW\Sigma W}^{T}\mathbf{J}^{T}$ (3)
Subsequently, an alpha-blend rendering technique, similar to that used in NeRF [36], is applied. This is formulated as:
$\mathbf{C}_{\text{color}}=\sum_{i\in N}c_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{i})$ (4)
Here, $c_{i}$ represents the color of each point, and the density $\alpha_{i}$ is determined by the product of a 2D Gaussian with covariance $\Sigma^{\prime}$ and a learned per-point opacity. The color is defined using spherical harmonics (SH) coefficients as described in [10, 27].
In summary, the basic 3D Gaussian Splatting approach characterizes each Gaussian point with the following attributes: (1) a 3D position $\mathcal{X}\in\mathbb{R}^{3}$, (2) a color defined by SH coefficients $c\in\mathbb{R}^{k}$ (where $k$ denotes the dimensionality of the SH basis), (3) a rotation represented by a quaternion $r\in\mathbb{R}^{4}$, (4) a scaling factor $s\in\mathbb{R}^{3}_{+}$, and (5) an opacity $\alpha\in[0,1]$.
3.2 Overview of the proposed GaussianOcc
The overview of the GaussianOcc is illustrated in Figure 2. The key contributions are the introduction of cross-view Gaussian splatting for scale-aware training and voxel grid Gaussian splatting for faster rendering. Unlike the original Gaussian splatting used for scene-specific 3D reconstruction, we explore a setting where Gaussian attributes are well-aligned in both 2D and 3D grids. This approach models the 3D scene in the 2D image plane as a depth map and in 3D grid space as a voxel (occupancy) format.
3.3 Scale-ware training by Gaussian Splatting
Scale from spatial camera rig: Similar to the previous work [16, 44], the scale information is from the surround camera rig. Specifically, the real-world scale could be obtained by leveraging camera extrinsic matrices which is to use spatial photometric loss in the overlap region between two adjacent views, i.e., warping ${I}^{i}_{t}$ to ${I}^{j}_{t}$:
$p^{i\rightarrow j}_{t}=K^{j}(T^{j})^{-1}T^{i}D_{t}^{i}(K^{i})^{-1}p^{i}_{t}$ (5)
where $K^{i},T^{i}$ are the intrinsic and extrinsic matrices of $i$th camera, $D_{t}^{i}$ is the predict depth map of $i$th camera, $p_{t}$ is the corresponding pixel during the warping. The warping operation is achieved by bilinear interpolation with the corresponding $p^{i\rightarrow j}_{t}$. However, as pointed out by [44], the mapping $p^{i\rightarrow j}_{t}$ is easy to be out of the image bounds with the weak supervision signal in such a small overlap region. Our practice also verifies this limitation. To provide a stronger supervision signal, [44] proposes to facilitate the Structure-from-Motion (SFM) to extract sparse depth information for the scale-aware training, which is time-consuming and not straightforward. Differently, we propose to use the Gaussian splatting for scale-aware projection to achieve the cross-view stereo constraint.
Gaussian Splatting for Projection (GSP): As illustrated in Figure 2, we use a depth network to predict the Gaussian attributes in 2D grid space, including the depth map, scale map, and rotation map. For each adjacent view, we first calculate the mask in the overlap region, then mask out one side of these overlap regions. Due to the presence of the other side’s overlap region, the unprojected 3D scene remains complete, which is critical for providing scale training, as indicated in the experiment section. We then perform splatting rendering on the adjacent views to obtain the rendered image. If the depth map is accurately learned, the rendered image should resemble the original images, providing the necessary scale information for the entire pipeline.
The process of acquiring the overlap mask is shown in Figure 3. The acquisition of the overlap mask is based on volume rendering. We densely sample points along the ray in one view, and a pixel is considered part of the overlap region if more than one sampled 3D point falls within the adjacent view. Note that in the DDAD dataset [14], we exclude regions with self-occlusion (such as parts of the vehicle body). Finally, we apply an erosion operation from Open CV library [2] to the mask for purification.
3.4 Fast rendering by Gaussian Splatting
Inefficient performance in volume rendering: For 2D supervision (semantic and depth maps), previous methods [12, 54, 24, 37] employed volume rendering based on dense sampling. While the final 3D voxel representation for modeling the 3D scene is much quicker than the original implicit representation (MLPs) [36], it remains time-consuming, particularly when incorporating semantic map rendering. For example, in OccNeRF [54], the number of sampled points at a resolution of $180\times 320$ is 108,735,066. However, the target optimized points correspond to the vertices in the 3D voxel grid, which total $320\times 320\times 24=2,457,600$. This redundancy in densely sampled points aids optimization with volume rendering but is highly inefficient.
Gaussian Splatting from Voxel (GSV): As analyzed above, the target optimized points are the vertices in the 3D voxel grid, prompting us to consider directly optimizing these vertices. Therefore, we investigate the use of Gaussian splatting to further accelerate the rendering process. For each vertex in the 3D voxel grid, we obtain the necessary attributes for Gaussian splatting, such as 3D position, semantic information, and opacity. Consequently, we propose treating each vertex as a 3D Gaussian. Since these vertices are arranged in the 3D voxel space, we can model the 3D scene by optimizing the related Gaussian attributes.
3.5 Loss function
We formulate the the total loss function in:
$\mathcal{L}_{total}=\mathcal{L}_{temporal}+\mathcal{L}_{cross}+\mathcal{L}_{semantic}+\mathcal{L}_{tv},$ (6)
$(\mathcal{L}_{temporal},\mathcal{L}_{cross})=\beta\frac{1-\operatorname{SSIM}\left(I_{t},\hat{I}_{t}\right)}{2}+(1-\beta)\left\|I_{t}-\hat{I}_{t}\right\|,$ (7)
where the terms $I_{t}$ and $\hat{I}_{t}$ refer to the target image and the corresponding synthesized image, respectively. Note that $\hat{I}_{t}$ in the temporal-view photometric loss, $\mathcal{L}_{temporal}$, is generated by projecting pixels from the source image using the coordinate index. In contrast, $\hat{I}_{t}$ in the cross-view photometric loss, $\mathcal{L}_{cross}$, is derived from our proposed cross-view Gaussian splatting method. $\mathcal{L}_{semantic}$ is the 2D semantic loss. $\mathcal{L}_{tv}$ is the regularization loss for the whole voxel space. The balancing parameter $\beta$ is set to 0.85 to balance the two loss terms.
Figure 3: Overlap mask in nuScenes [3] and DDAD [14].
Method
GT Occ.
GT Pose
mIoU mIoU*
$\blacksquare$ barrier
$\blacksquare$ bicycle
$\blacksquare$ bus
$\blacksquare$ car
$\blacksquare$ const. veh.
$\blacksquare$ motorcycle
$\blacksquare$ pedestrian
$\blacksquare$ traffic cone
$\blacksquare$ trailer
$\blacksquare$ truck
$\blacksquare$ drive. suf.
$\blacksquare$ sidewalk
$\blacksquare$ terrain
$\blacksquare$ manmade
$\blacksquare$ vegetation
MonoScene [4] $\times$ 6.33 6.06 7.23 4.26 4.93 9.38 5.67 3.98 3.01 5.90 4.45 7.17 14.91 7.92 7.43 1.01 7.65
BEVDet  [21] $\times$ 20.03 19.38 30.31 0.23 32.26 34.47 12.97 10.34 10.36 6.26 8.93 23.65 52.27 26.06 22.31 15.04 15.10
BEVFormer [30] $\times$ 24.64 23.67 38.79 9.98 34.41 41.09 13.24 16.50 18.15 17.83 18.66 27.70 48.95 29.08 25.38 15.41 14.46
OccFormer [56] $\times$ 22.39 21.93 30.29 12.32 34.40 39.17 14.44 16.45 17.22 9.27 13.90 26.36 50.99 34.66 22.73 6.76 6.97
RenderOcc [37] $\times$ 24.53 23.93 27.56 14.36 19.91 20.56 11.96 12.42 12.14 14.34 20.81 18.94 68.85 42.01 43.94 17.36 22.61
TPVFormer [23] $\times$ 28.69 27.83 38.90 13.67 40.78 45.90 17.23 19.99 18.85 14.30 26.69 34.17 55.65 37.55 30.70 19.40 16.78
CTF-Occ [41] $\times$ 29.54 28.53 39.33 20.56 38.29 42.24 16.93 24.52 22.72 21.05 22.98 31.11 53.33 37.98 33.23 20.79 18.00
SimpleOcc [12] $\times$ 7.99 7.05 0.67 1.18 3.21 7.63 1.02 0.26 1.80 0.26 1.07 2.81 40.44 18.30 17.01 13.42 10.84
SelfOcc [24] $\times$ 10.54 9.30 0.15 0.66 5.46 12.54 0.00 0.80 2.10 0.00 0.00 8.25 55.49 26.30 26.54 14.22 5.60
OccNeRF [54] $\times$ 10.81 9.54 0.83 0.82 5.13 12.49 3.50 0.23 3.10 1.84 0.52 3.90 52.62 20.81 24.75 18.45 13.19
GaussianOcc $\times$ $\times$ 11.26 9.94 1.79 5.82 14.58 13.55 1.30 2.82 7.95 9.76 0.56 9.61 44.59 20.10 17.58 8.61 10.29
Table 1: 3D Occupancy prediction performance on the Occ3D-nuScenes dataset in mIoU metric. Since ‘other’ and ‘other flat’ classes are the invalid prompts for open-vocabulary models, we also calculate ‘mIoU’ as the result ignoring the classes that do not consider these two classes during evaluation, while ‘mIoU*’ is the original result. GT Occ. means using the ground truth occupancy label for the supervision. GT Pose is the ground truth pose from the sensor for self-supervised geometry learning.
Method
GT Occ.
GT Pose
Backbone Input Size Epoch RayIoU RayIoU1m, 2m, 4m mIoU* FPS
[…]
Table 2: 3D Occupancy prediction performance on the Occ3D-nuScenes dataset in RayIoU metric. GT Occ. means using the ground truth occupancy label for the supervision. GT Pose is the ground truth pose from the sensor for self-supervised geometry learning. “8f” and “16f” mean fusing temporal information from 8 or 16 frames. mIoU* is the mean Intersection over Union for all categories. FPS means frame per second for each method, which is measured on a Tesla A100 GPU.
Method GT pose Occ. Abs Rel Sq Rel RMSE RMSE log $\delta<1.25$ $\delta<1.25^{2}$ $\delta<1.25^{3}$
nuScenes [3]
FSM [16] $\times$ $\times$ 0.297 - - - - - -
FSM* [16] $\times$ $\times$ 0.319 7.534 7.860 0.362 0.716 0.874 0.931
SurroundDepth [44] $\times$ $\times$ 0.280 4.401 7.467 0.364 0.661 0.844 0.917
SA-FSM [50] $\times$ $\times$ 0.272 4.706 7.391 0.355 0.689 0.868 0.929
VFF [28] $\times$ $\times$ 0.289 5.718 7.551 0.348 0.709 0.876 0.932
R3D3 [38] $\times$ $\times$ 0.253 4.759 7.150 - 0.729 - -
GaussianOcc $\ddagger$ $\times$ $\times$ 0.258 5.733 7.222 0.343 0.753 0.888 0.934
SimpleOcc [12] 0.224 3.383 7.165 0.333 0.753 0.877 0.930
OccNeRF [54] 0.202 2.883 6.697 0.319 0.768 0.882 0.931
SelfOcc [24] 0.215 2.743 6.706 0.316 0.753 0.875 0.932
GaussianOcc $\times$ 0.211 3.115 7.131 0.326 0.762 0.878 0.931
DDAD [14]
FSM*  [16] $\times$ $\times$ 0.228 4.409 13.433 0.342 0.687 0.870 0.932
VFF [28] $\times$ $\times$ 0.218 3.660 13.327 0.339 0.674 0.862 0.932
SurroundDepth [44] $\times$ $\times$ 0.208 3.371 12.977 0.330 0.693 0.871 0.934
SA-FSM [50] $\times$ $\times$ 0.187 3.093 12.578 0.311 0.731 0.891 0.945
R3D3 [38] $\times$ $\times$ 0.162 3.019 11.408 - 0.811 - -
GaussianOcc $\ddagger$ $\times$ $\times$ 0.212 3.556 12.564 0.320 0.701 0.888 0.944
GaussianOcc $\times$ 0.228 3.854 14.326 0.357 0.660 0.853 0.922
Table 3: Comparisons for self-supervised multi-camera depth estimation on the nuScenes  [3] and DDAD datasets  [14]. The results are averaged over all views without median scaling at test time. ‘FSM*’ is the reproduced result in [28]. GaussianOcc $\ddagger$ represents the depth estimation result from Stage 1. Occ. represents the ability of the method to predict the 3D occupancy.
Figure 4: The visualization of the render depth map and 3D occupancy prediction on nuScenes and DDAD dataset.
4 Experiment
4.1 Tasks, Datasets, and Metric
nuScenes: We perform 3D occupancy estimation and depth estimation using the nuScenes dataset [3]. For 3D occupancy estimation, we utilize annotations from Occ3D-nuScenes [41], which includes 600 scenes for training and 150 scenes for evaluation. Each sample has a perception range of [-40m, -40m, -1m, 40m, 40m, 5.4m] with a voxel size of 0.4m. For fair comparison, we use the 2D semantic map provided by OccNeRF [54]. We measure 3D occupancy estimation performance using the mean Intersection over Union (mIoU) metric. For depth estimation, we set the perception range in [-80m, -80m, -1m, 80m, 80m, 6m], while we clamp the ground truth to a range of 0.1m to 80m for evaluation, consistent with OccNeRF and SurroundDepth [44]. We evaluate depth maps using error metrics (Abs Rel, Sq Rel, RMSE, RMSE log) and threshold accuracy metrics ($\delta$).
DDAD: We conduct 3D occupancy estimation and depth estimation using the DDAD dataset [14]. Due to the lack of ground truth labels, we only present qualitative results for the fully self-supervised 3D occupancy estimation setting. Following the procedure used with the nuScenes dataset, we obtain the 2D semantic labels as per the OccNeRF pipeline. For depth estimation, we also set the perception range in [-80m, -80m, -1m, 80m, 80m, 6m], but clamp the ground truth to a range of 0.1m to 200m for evaluation, consistent with SurroundDepth [44].
Scale aware training in [44] Scale aware training by ours RMSE
Spatial loss SFM loss GS loss Mask Erode Refine Abs Rel Sq Rel RMSE log $\delta<1.25$ $\delta<1.25^{2}$ $\delta<1.25^{3}$
✓* 0.970 13.702 17.931 3.654 0.000 0.000 0.001
0.921 30.393 14.444 0.808 0.057 0.187 0.667
✓* ✓* 0.280 4.401 7.467 0.364 0.661 0.844 0.917
0.672 25.405 11.999 0.568 0.419 0.808 0.878
0.798 11.571 15.251 1.472 0.006 0.015 0.028
0.293 7.127 7.536 0.376 0.743 0.876 0.923
0.281 6.986 7.347 0.354 0.766 0.885 0.929
0.258 5.733 7.222 0.343 0.753 0.888 0.934
Table 4: Ablation study for scale-aware depth estimation on the Nuscene dataset [15]. The scores are averaged over all views without median-scaling at test time. ✓* means the result from the original paper and ✓means the result from our implementation in New-CRFs network [52]. Spatial loss denotes the spatial context loss used in [44] and SFM loss denotes the sparse depth pretrain from the off-line Structure-from-Motion process. GS loss means using the spatial context constraint by our proposed Gaussian splatting for projection. Mask represents using the mask-out strategy before the unprojection. Erode means use the erode process to the binary overlap mask and Refine is the refinement of the depth estimation network by fixing the 6D pose net.
4.2 Implementation details
Network details: For U-Net architecture, we adopt the New-CRFs [52] as the depth estimation network, which is based on the Swin Transformer [34]. The 6D pose network is the same as that used in SurroundDepth [44]. For the 2D to 3D lifting, we follow the approach used in SimpleOcc [12]. In depth estimation benchmark, we use the network proposed by SimpleOcc, where the final output size is 256×256×16. In our Gaussian splatting setting, we further upsample the final output to 512×512×32 for improved performance, which only requires minimal additional computational cost. For occupancy estimation, we use the same network as OccNeRF [54] to ensure a fair comparison.
Training details: We propose a two-stage training process for fully self-supervised 3D occupancy estimation. In the first stage, we jointly train the depth estimation network and the 6D pose network. In the second stage, we train the 3D occupancy network without optimizing the 6D pose network. During the first stage, we train the models for 8 epochs on the nuScenes dataset and 12 epochs on the DDAD dataset. In the second stage, we train the models for 12 epochs on both the nuScenes and DDAD datasets. The optimizer and learning rate adjustment strategy are the same as those used in SimpleOcc [12] and OccNeRF [54].
4.3 Main results
3D occupancy estimation in mIoU metric: As shown in Table 1, the proposed GaussianOcc achieves the best performance compared with other self-supervised methods, without the need for ground truth occupancy labels and ground truth 6D poses for training. It’s worth noting that SelfOcc [24] differs slightly from methods like SimpleOcc, OccNeRF, and our approach. Unlike these methods, SelfOcc does not predict semantic information directly in the 3D occupancy space. Instead, it attaches 2D semantic information to the 3D occupied voxels, where the 2D semantic data is sourced from a third-party open-vocabulary model [55].
3D occupancy estimation in RayIoU metric: We also compare our method using a newly proposed metric, RayIoU, introduced by the recent work [32]. The RayIoU is a ray-based evaluation metric to solve the inconsistency penalty along the depth axis raised in traditional voxel-level mIoU criteria. As shown in Table 2, our approach also outperforms OccNeRF [54]. It’s important to note that the FPS is calculated excluding rendering time. Since GaussianOcc and OccNeRF utilize the same network architecture, they share the same inference time when the rendering process is not taken into account.
3D occupancy estimation in DDAD: Thanks to our fully self-supervised learning approach, we are able to perform 3D occupancy estimation on a similar dataset, DDAD [14]. To the best of our knowledge, our work is the first to introduce 3D occupancy estimation on this dataset, although we cannot provide quantitative results due to the lack of 3D annotations. We present the visualization results in Figure 4 and more results could be found in the supplement material.
Depth estimation: We present a comparison of depth estimation results in Table 3 for both the nuScenes and DDAD datasets. In the first stage of monocular depth estimation, GaussianOcc $\ddagger$ achieves top performance on the nuScenes dataset and delivers competitive results on the DDAD dataset. It is important to note that methods such as SurroundDepth [44] and SA-FSM [50] rely on third-party sparse point cloud extraction. Additionally, R3D3 [38] is a temporal offline refinement method that requires multi-frame optimization, as discussed in [24]. In the second stage, which involves depth estimation from rendering, our method also achieves competitive results compared to those trained with ground truth poses. Besides, we observe that the rendered depth in stage two outperforms the depth results from stage 1 on the nuScenes dataset, whereas the opposite is true for the DDAD dataset. This discrepancy might be attributed to differences in perception range—80 meters in nuScenes versus 200 meters in DDAD.
4.4 Ablation Study
Scale aware training: To demonstrate the effectiveness of the proposed scale-aware training using Gaussian Splatting, we conducted experiments comparing our method with existing approaches that use SFM loss for training [44]. As indicated in Table 4, we implemented scale-aware training from [44] on our network [52], but the results were suboptimal, and our proposed method outperformed the results reported in the original paper. In our method, we observed that a naive implementation of Gaussian splatting without a masking strategy is ineffective because it lacks cross-view rendering, which is essential for applying cross-view constraints. To enhance performance, we introduced an erosion operation on the binary mask to purify it, excluding regions that may fall outside the overlap area. This step ensures better alignment during training. Finally, we refined the depth estimation by fixing the 6D pose network and disabling the cross-view loss. This refinement helps reduce artifacts at the edges of the overlap region. We present the cross-view splatting render result in Figure 5, where the rendering result in overlap regions is reasonable in the correct depth estimation from cross view.
6D pose and training strategy: Building on the success of scale-aware training, we compared the depth estimation performance of our proposed two-stage training method with that of models trained using ground truth (GT) pose. As shown in Table 5, our two-stage approach delivers competitive results when compared to training with GT pose, indicating that the predicted pose is of high quality. Additionally, the GT pose provided by the nuScenes dataset is limited by sensor inaccuracies, particularly a lack of vertical movement. In contrast, the learned 6D pose in our method does not suffer from this limitation. Furthermore, in our one-stage experiment, we found that this approach was not effective. The cross-view loss in the 3D voxel space led to local optimization, which failed to generalize predictions to non-overlapping regions as indicated in Figure 6. However, this issue did not arise in the depth maps generated by the 2D decoder in stage one.
Volume rendering and Splatting rendering: In Table 6, we examine the performance of volume rendering (VR) and splatting rendering (SR) in the voxel space. For the scale of each 3D Gaussian, we set a uniform scale $s$ for all Gaussians, considering the well-arranged positions of vertices within the voxel grid. Since the scale $s$ is consistent across all three dimensions, we set the rotation $\mathbf{R}$ as the identity matrix. We first explored the optimal scale and found that a scale of 0.1 produced the best results. Notably, the performance of VR and SR was very close when using this scale. The render depth maps are shown in Figure 6. The depth map by the volume rendering is smoother thanks to dense sampling and the depth maps with splatting tendering have the graininess effect, especially at the small scale factor.
Pose and render types ablation study in occupancy task: In Table 7, we conduct an ablation study to evaluate the impact of using ground truth (GT) versus learned poses in the occupancy task, with training conducted using different rendering methods. The results show that splatting rendering achieves superior performance in both occupancy metrics (mIoU) and depth metrics. One observation is that depth estimation results from 3D semantic learning with volume rendering are significantly worse than those in Table 3 and the proposed splatting rendering method maintains consistent performance, which indicates the splatting rendering in voxel grid may contribute better 3D geometry.
Rendering efficiency analysis: Finally, we analyze rendering efficiency in Table 8. The results indicate that volume rendering consumes 5 times more rendering time compared to splatting rendering, and as the resolution increases, both rendering time and GPU consumption rise significantly. In contrast, splatting rendering shows no significant increase in computational cost with higher resolutions, highlighting its efficiency and scalability.
Pose type Abs Rel Sq Rel RMSE $\delta<1.25$
GT pose 0.214 3.362 7.127 0.771
Two stages 0.211 3.115 7.131 0.762
One stage 0.946 17.008 16.397 0.103
Table 5: Comparison of the ground truth pose (GT pose) and our learned pose (Two stages) on the depth estimation task [15]. We also conducted the experiment that directly uses the cross-view loss to render a depth map (One stage).
Render type Abs Rel Sq Rel RMSE $\delta<1.25$
VR 0.215 3.508 7.113 0.775
SR ($s$ = 0.05) 0.223 3.694 7.246 0.761
SR ($s$ = 0.1) 0.217 3.504 7.152 0.770
SR ($s$ = 0.15) 0.217 3.406 7.204 0.763
Table 6: Comparison of the render result between the volume rendering (VR) [54] and splatting rendering (SR, Ours) on depth estimation task [15]. The coefficient $s$ means the scale of the 3D Gaussian. We use GT pose for this set of the ablation study.
Pose type mIoU Abs Rel Sq Rel $\delta<1.25$
GT pose (VR) 10.81 0.456 12.682 0.704
GT pose (SR) 11.30 0.225 4.339 0.787
Learned pose (VR) 11.19 0.506 15.577 0.684
Learned pose (SR) 11.43 0.219 3.625 0.778
Table 7: Comparison of ground truth pose (GT Pose) and our learned pose (Two stages) on 3D occupancy estimation task [15] in volume rendering (VR) and splatting rendering (SR).
Render Render resolution and time (s) Training
type 180 $\times$ 320 240 $\times$ 520 360 $\times$ 640 time (h)
VR $\approx$ 0.85 $\approx$ 1.57 N/A $\approx$ 2.68
SR $\approx$ 0.17 $\approx$ 0.17 $\approx$ 0.17 $\approx$ 1
Table 8: Comparison of rendering efficiency between volume rendering (VR) [54] and splatting rendering (SR, Ours) on 3D occupancy estimation task [15]. The render time is calculated from surround 6 images. ”N/A” indicates out-of-memory errors running in NVIDIA A 100 (40 GB). Training time is averaged per epoch.
Figure 5: The visualization for the depth map, splatting render image and depth map from cross view in stage 1.
Figure 6: The comparison for the depth map in the different setting, corresponding to the training strategy in Table 5 and rendering type in Table 6.
5 Conclusion
In this paper, we introduce GaussianOcc, a fully self-supervised and efficient approach for 3D occupancy estimation. Through the carefully designed cross-view splatting rendering, we can accurately learn the real scale in depth and the 6D pose net, enabling effective self-supervised 3D occupancy learning. Additionally, the proposed Gaussian splatting in voxel grids outperforms volume rendering in 3D occupancy estimation while reducing computational cost. We validate our method on the nuScenes and DDAD datasets, demonstrating its strong generalization ability to surround view scenes.
Appendix
Appendix A More implementation details
The detailed parameter setting in depth estimation network [52]. During the joint depth and 6D pose training in stage 1, we predict the 3D Gaussian parameters alongside the 2D depth map. Since the Gaussian parameters are well-arranged in the 2D image plane prior to unprojection, we maintain equal scaling across all three dimensions of each 3D Gaussian and constrain the maximum scale to 0.02. Given that the scale $s$ is uniform across all dimensions, we set the rotation matrix $\mathbf{R}$ to the identity matrix. Additionally, we assign an opacity value of 1 to each 3D Gaussian, ensuring that every 2D depth value corresponds to a valid point in 3D space.
The detailed parameter setting in voxel grid splatting rendering. We chose a fixed scale for each grid vertex to ensure a well-arranged structure that accurately models the 3D space. If we were to use a learnable scale, it could lead to a situation where the scale is small but the opacity is large, which might not be captured in the rendered depth map and semantic map but could still impact the 3D occupancy result. Therefore, using a fixed scale is sufficient for optimization, as demonstrated in the results presented in the main paper.
[…]
Detailed definition of depth map metric. Following the depth estimation task [44], we report the depth map evaluation with the following metrics,
[…]
where $M$ is the valid pixel, $\hat{d}$ is the ground truth depth and $d^{*}$ is the predicted depth.
Appendix B More experiment results
More visualization. We provide more visualization for nuScenes dataset in Figure A and DDAD dataset in Figure B. Please check the attached videos for sequence visualization.
Appendix C Limitation and future work
The proposed method achieves reasonable predictions in most scenes; however, we observe that some cases still present challenges, as shown in Figure C. Specifically, in the DDAD dataset, incorrect predictions occur in the back camera in certain situations as marked with the red circle, where the drivable surface is mistakenly projected into the car due to extensive self-occlusion. Notably, this issue is absent in the nuScenes dataset, which has less self-occlusion. We believe that this problem could be mitigated with better 2D semantic maps for supervision, which warrants further investigation. Additionally, we plan to explore the use of splatting rendering for 3D occupancy flow in future work.
Figure A: The visualization of the render depth map and 3D occupancy prediction on nuScenes dataset.
Figure B: The visualization of the render depth map and 3D occupancy prediction on DDAD dataset
Figure C: Some wrong predictions due to the large self-occlusion on DDAD dataset.